# Intrinsic Heisenberg-type lower bounds on spacelike hypersurfaces in general relativity

### v2

@@ 3.2 Barta-type improvement under weak mean-convexity
Under Assumption 2.1, the identity $d=r-\rho$ reduces the boundary-distance Hardy hypothesis to the radial condition used in the Barta argument. Thus the improvement from the universal Hardy baseline (19) to the Barta-type estimate (23) comes from the vector-field method and choice of $X$, not from a stronger geometric assumption. This improvement is scale-invariant and uses only the sign of $\Delta\rho$, hence it is robust, geometry- and foliation-independent within the present framework. In minimax sense and three dimensions, (23) is best possible under exactly the hypotheses of Corollary 3.7; any further strengthening requires additional information (e.g. curvature/comparison hypotheses), cf. [5, 6] or model-specific refinement [10].
4 Discussion, scope, and extensions
The intrinsic lower bound of Theorem 3.1 shows that momentum uncertainty on a spacelike slice is governed solely by the Dirichlet spectrum of the induced Riemannian metric. The estimate is therefore insensitive to the lapse, the shift, and the extrinsic curvature, and it remains stable under changes of foliation. In combination with the exact variance decomposition, the kinetic-energy window (18) separates two contributions to $\sigma_{p}$: the geometric cost encoded in the modulus $|\psi|$ and the fluctuations of the phase gradient. The lower endpoint is controlled by the ground-state Dirichlet mode of $\Delta_{h}$, while the upper endpoint is attained precisely when the mean momentum vanishes. In this sense the result provides a coordinate-independent bridge between spectral geometry and an operational notion of uncertainty.
Under Assumption 2.1 and weak mean-convexity of the geodesic boundary, the sharp boundary-distance Hardy inequality yields the universal product estimate (19). This floor is scale-invariant in $r$ and depends only on the boundary distance $d(x)$; neither interior curvature nor extrinsic spacetime data enter. Consequently, even without detailed geometric information one obtains a robust, geometry-free lower bound on $\sigma_{p}$ that is available on sufficiently small balls (below the convexity radius) and in many natural situations. The constant is optimal and never attained, mirroring classical features of Hardy-type inequalities.
Beyond the Hardy floor a factor-$\pi$ improvement is obtained via the vector-field version of Barta’s method [5]. Together with Theorem 3.1, this yields a strictly stronger, model-agnostic control on weakly mean-convex balls, and it can only be further sharpened wherever eigenvalue comparison and curvature-sensitive techniques are available [6, 7].
Classical eigenvalue comparison theorems under sectional or Ricci curvature bounds (Cheng-type inequalities) yield effective upper and lower bounds for the first Dirichlet eigenvalue of geodesic balls; see [6, 7]. In the small-radius regime, asymptotic expansions quantify the departure of $\lambda_{1}$ from its Euclidean value in terms of curvature at the center point [8]. On spherically symmetric manifolds, the radial structure can be exploited to compute or bound spectra explicitly [9]. Each of these inputs transfers directly to $\sigma_{p}$ via Theorem 3.1.
The Dirichlet framework enforces strict localization and expresses the variance purely through the bulk Dirichlet form, which makes spectral control transparent. On connected domains with Neumann data the ground state is constant and the spectral gap closes, so any induced lower bound degenerates. For Robin data a nonzero surface term persists in the variance, and the uncertainty can no longer be written purely in terms of the bulk form; any lower bound must then incorporate explicit boundary contributions. This clarifies why Dirichlet data are the natural modeling choice for the intrinsic statement proved here.
The argument is not restricted to scalar fields. Replacing $\Delta_{h}$ by the Dirac operator or the Hodge Laplacian yields analogous intrinsic lower bounds for spinors or differential forms, controlled by the corresponding first Dirichlet eigenvalues (up to operator-dependent constants). Foliation independence is retained because only the induced Riemannian data on the slice enter. Additional structures – such as gauge potentials or symmetry constraints – can be incorporated via standard variational methods together with diamagnetic or Hardy-type tools.
From a computational standpoint, standard finite-element or spectral discretizations of $\lambda_{1}$ make the estimate numerically actionable for concrete geometries (for instance, in numerical relativity). Physically, the window (18) clarifies near-saturation behavior: equality at the upper endpoint corresponds to vanishing mean momentum, whereas equality at the lower endpoint forces the frame derivatives $X_{a}(\arg\psi)$ to be almost everywhere constant. An explicit Euclidean model with ground-state modulus and affine phase exhibits saturation of the intrinsic bound with nonzero mean momentum.
Taken together, the universal Hardy/Barta baseline and curvature-sensitive refinements yield a two-tier structure: a geometry-free minimum available under weak mean-convexity and, when curvature information is present, sharper controls tailored to the ambient geometry. This template extends naturally to other strictly local observables, to anisotropic comparison results adapted to symmetry classes or foliations, and to measurement-theoretic formulations of quantum mechanics on curved backgrounds.

### v4

@@ 3.2 Barta-type improvement under weak mean-convexity
Context for the bounds (26) and (30). Under Assumption 2.1, the identity $d=r-\rho$ reduces the boundary-distance Hardy hypothesis to the radial condition used in the Barta argument. Thus the improvement from the universal Hardy baseline (26) to the Barta-type estimate (30) comes from the vector-field method and choice of $X$, not from a stronger geometric assumption. This improvement is scale-invariant and uses only the sign of $\Delta\rho$, hence it is robust, geometry- and foliation-independent within the present framework. In minimax sense and three dimensions (see, e.g., the discussion in [18]), (30) is best possible under exactly the hypotheses of Corollary 3.7; any further strengthening requires additional information (e.g. curvature/comparison hypotheses), cf. [18, 19] or curvature-dependent refinement [25].
4 Examples and applications
An illuminating perspective on the geometric potential $V$ arises when the Riemannian manifold $(\Sigma,h)$ is constructed from a Lie algebra. Instead of deriving the frame from a metric, one specifies the commutation relations of an orthonormal frame $\{X_{a}\}_{a=1}^{n}$,
$X_{a},X_{b}]=\sum_{c=1}^{n}C_{ab}^{c}X_{c},$ (33)
where $C_{ab}^{c}$ are the structure constants satisfying the Jacobi identity. This construction yields a Lie group (or a quotient thereof) equipped with a left-invariant metric $h(X_{a},X_{b})=\delta_{ab}$. In this algebraic setting, the divergence of the frame fields is constant and determined by the trace of the structure constants (adjoint representation),
$f_{a}=\mathrm{div}_{h}X_{a}=-\sum_{b=1}^{n}C_{ab}^{b}.$ (34)
Since $f_{a}$ is constant, the derivative term in (6) vanishes, and the geometric potential becomes a pure algebraic invariant:
$V=\frac{1}{4}\sum_{a=1}^{n}f_{a}^{2}\;>\;0.$ (35)
This classification creates a dichotomy between unimodular and non-unimodular geometries, which directly impacts the intrinsic momentum uncertainty.
Example 4.1
[Category A: Unimodular geometry / Heisenberg group]
Consider the 3-dimensional Heisenberg algebra $\mathfrak{h}_{3}$ generated by $\{X_{1},X_{2},X_{3}\}$ with the non-vanishing commutator
$X_{1},X_{2}]=X_{3}.$ (36)
A concrete realization of this algebra is provided by the Nil geometry, one of the eight Thurston model geometries. On the manifold $\Sigma\cong\mathbb{R}^{3}$ with coordinates $(x,y,z)$, the orthonormal frame generators are explicitly given by:
$X_{1}=\partial_{x}-\frac{1}{2}y\partial_{z},\quad X_{2}=\partial_{y}+\frac{1}{2}x\partial_{z},\quad X_{3}=\partial_{z}.$ (37)
The Riemannian metric $h$ that renders this frame orthonormal is the left-invariant metric
$ds^{2}_{h}=dx^{2}+dy^{2}+\left(dz+\frac{1}{2}(ydx-xdy)\right)^{2}.$ (38)
Returning to the algebraic perspective, the only non-zero structure constants are $C_{12}^{3}=-C_{21}^{3}=1$. Computing the divergences $f_{a}=-\sum_{b}C_{ab}^{b}$ yields $f_{1}=f_{2}=f_{3}=0$. Consequently, the algebra is unimodular, and the geometric potential vanishes identically, $V\equiv 0$.
Although the geometric potential $V$ vanishes identically, the geometry is non-flat: the scalar curvature is constant and negative, $\mathrm{Scal}=-\tfrac{1}{2}$. Moreover, for the maximally symmetric left-invariant Riemannian Nil metric considered here, the injectivity radius is finite and spatially constant, $\operatorname{inj}_{\Sigma}(p)\equiv 2\pi$ for all $p\in\Sigma$. Consequently, the strict localization assumption (Assumption 2.1) imposes a specific upper bound on the size of the geodesic ball in this geometry, requiring
$0<r<2\pi\,.$
The value of the injectivity radius and its point-independence follow from the explicit description of the geodesic flow, conjugate and cut loci for left-invariant Riemannian metrics on the Heisenberg groups in Biggs and Nagy [32].
Example 4.2
[Category B: 3D Hyperbolic Space $\mathbb{H}^{3}$]
An analytically tractable example of a non-unimodular geometry is the three-dimensional hyperbolic space $\mathbb{H}^{3}$ with curvature $\kappa=-1$. We model $\mathbb{H}^{3}$ as the group manifold generated by two translations $\{X_{1},X_{2}\}$ and one dilation $X_{3}$, satisfying
$X_{3},X_{1}]=X_{1},\qquad[X_{3},X_{2}]=X_{2},\qquad[X_{1},X_{2}]=0.$
The non-vanishing structure constants are $C_{31}^{1}=C_{32}^{2}=1$. Summing over the adjoint indices yields the divergence vector of the frame: $f_{3}=-(C_{31}^{1}+C_{32}^{2})=-2$ and $f_{1}=f_{2}=0.$ Inserting this into the definition of the geometric potential (6) gives the constant algebraic shift
$V=\frac{1}{4}f_{3}^{2}=1\,.$
Conversely, consider the Dirichlet problem on a geodesic ball of radius $R$ in $\mathbb{H}^{3}$. The radial Laplacian $\Delta_{r}=\partial_{r}^{2}+2\coth(r)\partial_{r}$ transforms via $u(r)=\chi(r)/\sinh(r)$ into the Schrödinger operator $\partial_{r}^{2}+1$. The ground state $\chi(r)=\sin(\pi r/R)$ corresponds to the eigenvalue
$\lambda_{1}(\mathbb{H}^{3},R)=\frac{\pi^{2}}{R^{2}}+1.$
The algebraic potential $V=1$ exactly matches the spectral gap $\lim_{R\to\infty}\lambda_{1}=1$. The intrinsic Heisenberg bound thus takes the explicit form [25]
$\sigma_{p}(\psi)^{2}\;\geq\;\hbar^{2}\left(\frac{\pi^{2}}{R^{2}}+1\right),$
demonstrating how the non-unimodular algebra enforces a strictly positive lower bound on momentum uncertainty even in the infinite-volume limit. Using the variance decomposition (7), we can isolate the naive variance measured by the momentum operators in this specific frame. For the ground state, we obtain:
$\mathrm{Var}^{\mathrm{naive}}_{\psi}(P)$ $\;=\;\hbar^{2}\frac{\,\pi^{2}}{R^{2}}.$
Strikingly, the naive variance recovers the exact Euclidean scaling $1/R^{2}$. The geometric potential $V$ acts as a filter that subtracts the spectral gap (the pure curvature energy) from the total intrinsic uncertainty, leaving behind the pure confinement energy associated with the ball size $R$. The injectivity radius is infinite, $\operatorname{inj}_{\Sigma}=\infty$, implying that the radius $r$ of the geodesic ball is unconstrained.
Example 4.3
[Non-constant curvature: “Witten’s Cigar”]
To demonstrate the scope of the intrinsic bound beyond homogeneous spaces, we consider the 2-dimensional manifold $\Sigma\cong\mathbb{R}^{2}$ equipped with the rotationally symmetric metric
$h\;=\;\mathrm{d}r^{2}+\tanh^{2}(r)\,\mathrm{d}\phi^{2},\qquad r\geq 0,\ \phi\in[0,2\pi).$ (39)
This geometry, known as Witten’s cigar or the Euclidean 2D black hole, interpolates smoothly between a spherical cap of positive Gaussian curvature $K=2$ at the origin and a flat cylinder ($K\to 0$) at infinity. A natural orthonormal frame $\{X_{a}\}$ is obtained by normalizing the coordinate basis:
$X_{1}=\partial_{r},\qquad X_{2}=\coth(r)\,\partial_{\phi}.$ (40)
Unlike in the Heisenberg or hyperbolic cases, the algebra of these fields is governed by position-dependent structure functions rather than constant coefficients. A direct computation of the Lie bracket yields
$X_{1},X_{2}]\;=\;-\frac{2}{\sinh(2r)}\,X_{2}.$ (41)
The non-vanishing structure function $C_{12}^{2}(r)=-2/\sinh(2r)$ encodes the inhomogeneity of the curvature. The divergences of the frame fields are $f_{1}=\mathrm{div}_{h}X_{1}=2/\sinh(2r)$ and $f_{2}=0$. Consequently, the geometric potential $V$ defined in (6) becomes a spatially varying field:
$V(r)\;=\;\frac{1}{2}\partial_{r}f_{1}+\frac{1}{4}f_{1}^{2}\;=\;\frac{1-2\cosh(2r)}{\sinh^{2}(2r)}.$ (42)
The potential (42) captures the global topology of the manifold:
  • As $r\to\infty$, the potential decays exponentially, $V(r)\to 0$, consistent with the asymptotic flatness of the cylinder.
  • As $r\to 0$, it diverges as $V(r)\sim-1/(4r^{2})$. This singular behavior is not pathological but physical: it exactly cancels the centrifugal barrier term coming from the Laplacian in polar coordinates, ensuring that the effective Hamiltonian is regular at the origin (where the geometry closes smoothly like $\mathbb{R}^{2}$).
Finally, the first Dirichlet eigenvalue $\lambda_{1}$ on a ball $B_{R}(0)$ is determined by the zeros of the Legendre function $P_{-1/2+i\nu}(\cosh 2R)$. Theorem 3.1 thus provides a strictly intrinsic lower bound $\sigma_{p}\geq\hbar\sqrt{\lambda_{1}}$ that automatically incorporates these non-trivial geometric features. For a detailed comparison between the intrinsic lower bound derived in Theorem 3.1 and classical spectral estimates in different radial regimes we consider the spectral asymptotics and sharpness of the bound:
  • Small-ball regime ($R\ll 1$). Near the origin, the geometry is dominated by the positive Gaussian curvature $K(0)=2$. The asymptotic expansion for small geodesic balls [23] yields
    $\lambda_{1}(B_{R})\;=\;\frac{j_{0,1}^{2}}{R^{2}}-\frac{K(0)}{3}+\mathcal{O}(R^{2}),$ (43)
    where $j_{0,1}\approx 2.405$ is the first zero of the Bessel function $J_{0}$. This confirms that the positive curvature at the tip acts as an attractive potential, lowering the momentum uncertainty relative to the flat Euclidean case.
  • Large-ball regime ($R\to\infty$). As the radius increases, the manifold transitions into a cylinder $\mathbb{R}\times S^{1}$. The radial Laplacian simplifies asymptotically to $\partial_{r}^{2}$, effectively reducing the problem to a one-dimensional particle in a box of length $R$. Consequently, the eigenvalue scales as
    $\lambda_{1}(B_{R})\;\sim\;\frac{\pi^{2}}{4R^{2}}\qquad(R\to\infty).$ (44)
    Remarkably, this asymptotic behavior coincides exactly with the universal Barta-type lower bound derived in Corollary 3.7 (with $\pi/2\approx 1.57$). This demonstrates that our coordinate-independent lower bound $\sigma_{p}r\geq\pi\hbar/2$ is asymptotically optimal for geometries that degenerate into lower-dimensional effective spaces, capturing the transition from 2D radial symmetry to 1D longitudinal confinement.
Injectivity Radius. Since the manifold is topologically $\mathbb{R}^{2}$ and the radial function $\tanh(r)$ is strictly monotonic, there are no conjugate points along radial geodesics from the origin. Thus, $\operatorname{inj}_{\Sigma}(0)=\infty$, allowing for arbitrarily large geodesic balls as discussed in the asymptotic analysis.
Example 4.4
[Injectivity radius at the Schwarzschild throat]
In general there is no simple closed-form expression for the injectivity-radius of a generic Riemannian $3$-manifold, and the maximal Schwarzschild slice is no exception. For the purposes of Assumption 2.1, however, one does not need an explicit formula for $\mathrm{inj}_{\Sigma}(x)$ at every point $x\in\Sigma$. The assumption only requires that the radius $r$ of the geodesic ball $B_{\Sigma}(p,r)$ be strictly smaller than $\mathrm{inj}_{\Sigma}(p)$ at the chosen center $p$. It is therefore natural to look for a point where the injectivity radius is smallest: any lower bound obtained from such a “worst case” automatically controls all admissible localization radii on the slice. On the time-symmetric Cauchy slice of the maximal Schwarzschild extension this role is played by the Einstein–Rosen throat [33].
More concretely, in the maximal Schwarzschild extension the time-symmetric Cauchy slice $(\Sigma,h)$ can be written in warped-product form
$h=d\ell^{2}+R(\ell)^{2}d\Omega^{2},$
where $\ell\in\mathbb{R}$ is a proper-distance coordinate and $R(\ell)$ is the area radius of the $2$-spheres $\{\ell=\mathrm{const}\}$. The minimal $2$-sphere (Einstein–Rosen throat / apparent horizon) is located at $\ell=\ell_{0}$ with $R(\ell_{0})=2M$ and $R^{\prime}(\ell_{0})=0$; the corresponding $2$-sphere is totally geodesic and carries the round metric of radius
$R_{\min}:=2M.$
Let $p\in\Sigma$ lie on this minimal $2$-sphere $S_{H}$. Geodesics in $(\Sigma,h)$ that are tangent to $S_{H}$ at $p$ remain in $S_{H}$ and coincide with great-circle geodesics on the round $2$-sphere of radius $R_{\min}$. These great circles have sectional curvature $1/R_{\min}^{2}$, and along them the first conjugate point occurs at distance $\pi R_{\min}$. Hence the conjugate radius at $p$ is
$\mathrm{conj}_{\Sigma}(p):=\pi R_{\min}=2\pi M.$
Moreover, the shortest nontrivial closed geodesics through $p$ are precisely these great circles, of length $2\pi R_{\min}=4\pi M$, so that half their length is again $2\pi M$. By Klingenberg’s injectivity-radius formula,
$\mathrm{inj}_{\Sigma}(p)=\min\Big\{\mathrm{conj}_{\Sigma}(p),\tfrac{1}{2}\ell_{p}\Big\}=2\pi M,$
where $\ell_{p}$ denotes the length of the shortest nontrivial closed geodesic through $p$. Thus the injectivity radius at the Schwarzschild throat is finite and of order $M$. In particular, for any $p\in S_{H}$ the geodesic balls $B_{\Sigma}(p,r)$ satisfy Assumption 2.1 for every radius
$0<r<2\pi M,$
and the value $2\pi M$ can be regarded as a natural lower bound for the injectivity radius scale relevant to our intrinsic Heisenberg-type estimates. Since this quantity is derived exclusively from intrinsic metric data, it serves as a coordinate-independent lower bound for the localization scale.
Example 4.5
[Isotropic maximum-length model of GUP]
As a final example we consider the isotropic maximum-length (ML) model introduced in the context of the GUP [9, 10, 11, 12, 13]. Its kinematics is defined algebraically by deformed position-momentum commutation relations on $\mathbb{R}^{3}$ with Cartesian coordinates $x_{i}$,
$\hat{x}_{i},\hat{x}_{j}]=0,\qquad[\hat{x}_{i},\hat{P}_{j}]=i\hbar\,\frac{\delta_{ij}}{1-\alpha r^{2}},\qquad r^{2}:=x_{1}^{2}+x_{2}^{2}+x_{3}^{2},$ (45)
where $\alpha>0$ is a deformation parameter. In this algebraic formulation, the factor $(1-\alpha r^{2})^{-1}$ is the only modification of the usual flat-space commutator; the position operators remain mutually commuting. Applying a metric-reconstruction procedure with the general commutator rule (2), one reads off the orthonormal frame fields from the mixed commutator:
$X_{j}(x_{i})=\frac{\delta_{ij}}{1-\alpha r^{2}}\quad\Longrightarrow\quad X_{j}=\frac{1}{1-\alpha r^{2}}\,\frac{\partial}{\partial x_{j}}.$ (46)
Requiring $\{X_{j}\}$ to be orthonormal determines the inverse metric and hence the configuration-space metric of the ML model,
$g_{ij}(x)=(1-\alpha r^{2})^{2}\,\delta_{ij},$ (47)
so that the configuration space is the conformally flat Riemannian manifold $(\mathbb{R}^{3},g_{\mathrm{ML}})$ with conformal factor $1{-}\alpha r^{2}$. In this way, the “maximum-length” deformation of the flat-space commutator is reinterpreted as standard quantum mechanics on the curved manifold $(\mathbb{R}^{3},g_{\mathrm{ML}})$, with the entire modification encoded in the conformal factor of the metric, the induced symplectic structure, and the momentum commutator. The Ricci tensor of $g_{\mathrm{ML}}$ in Cartesian coordinates $(x_{1},x_{2},x_{3})$ is
$\operatorname{Ric}_{ij}(x)=\frac{8\alpha}{(\alpha r^{2}-1)^{2}}\bigl(\alpha x_{i}x_{j}-(\alpha r^{2}-1)\delta_{ij}\bigr),\qquad r^{2}=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$ (48)
and therefore the ML metric satisfies the uniform Ricci lower bound
$\operatorname{Ric}_{g_{\mathrm{ML}}}(X,X)\;\geq\;8\alpha\,g_{\mathrm{ML}}(X,X),\qquad\forall X\in T_{x}\mathbb{R}^{3},\;r(x)<\alpha^{-1/2}.$ (49)
In the standard notation $\operatorname{Ric}\geq(n-1)K\,g$ with $n=3$ this corresponds to $K=4\alpha$. This Ricci lower bound, together with nonnegative mean curvature of $\partial B_{r}$, allows one to apply Yang–Ling type inequalities [21, 22] to obtain
$\lambda^{\mathrm{ML}}_{1}\;\geq\;\frac{\pi^{2}}{4r^{2}}+4\alpha,$ (50)
where we have quoted Ling’s stronger bound [22]. By way of comparison, Barta’s inequality (Corollary 3.7) yields
$\lambda^{\mathrm{ML}}_{1}\;\geq\;\frac{\pi^{2}}{4r^{2}}\,.$ (51)
In particular, inequality (51) captures the universal $r^{-2}$-behavior of the first Dirichlet eigenvalue, whereas the refined estimate (50) improves this bound by an additional strictly positive, curvature-induced term $4\alpha$, which reflects the presence of an intrinsic spectral gap for the ML model. Moreover, for the ML metric the injectivity radius depends on the base point $p\in\Sigma$ and satisfies the bounds
$0\;\leq\;\mathrm{inj}_{\mathrm{ML}}(p)\;\leq\;\frac{2}{3\sqrt{\alpha}},\qquad p\in\Sigma.$ (52)
Hence the maximal admissible radius of geodesic balls centered at the origin is $r<\mathrm{inj}_{\mathrm{ML}}(0)=2/(3\sqrt{\alpha})$, while the injectivity radius tends to zero as $p$ approaches the Euclidean boundary $r=1/\sqrt{\alpha}$ of the ML configuration space.
Taken together, the comparison inequalities for $\lambda^{\mathrm{ML}}_{1}$ and the injectivity-radius estimate (52) provide a global control of the spectrum in the maximum-length geometry: for every base point ($p$) and every admissible localization radius ($r<\mathrm{inj}^{\mathrm{ML}}(p)$) the first Dirichlet eigenvalue is bounded from below by an $r^{-2}$- term, and the Yang-Ling refinement adds a strictly positive curvature-induced contribution. From the perspective of the GUP phenomenology, however, the most relevant regime is precisely the opposite one, in which the localization scale is small compared to the curvature scale set by ($\alpha^{-1/2}$), so that the deformation away from flat space is weak.
It is therefore natural to complement these global comparison bounds by a local small-ball analysis around the origin in which the metric $g_{\mathrm{ML}}$ may be treated as a mild conformal perturbation of the Euclidean metric. Specializing to geodesic balls centered at the origin with radius ($r$) satisfying $\alpha r^{2}\ll 1$ and hence $r<\mathrm{inj}_{\mathrm{ML}}(0))$, the Dirichlet problem becomes accessible to a straightforward radial perturbation calculation. In this perturbative regime one finds the asymptotic expansion
$\lambda^{\mathrm{ML}}_{1}(r)=\frac{\pi^{2}}{r^{2}}-4\alpha+O(\alpha^{2}r^{2}),$ (53)
so that the leading $r^{-2}$ behaviour coincides with that of a geodesic ball in a three-sphere $S^{3}$ of constant sectional curvature $4\alpha$; the Ricci term $4\alpha$ is the first curvature correction.
5 Discussion and conclusion
The analysis carried out in this work revisits Heisenberg’s uncertainty principle in a preparation-based form that is adapted to curved spacetime. Instead of deforming the canonical commutation relations or postulating a specific generalized uncertainty principle, we have kept standard quantum mechanics on a fixed spacetime background and encoded the influence of gravity in the geometry of the localization region itself. Strict von Neumann–Lüders localization to a geodesic ball $B_{\Sigma}(p,r)$ on a spacelike hypersurface $(\Sigma,h)$ was modeled by Dirichlet data, and the resulting momentum uncertainty was described by a geometric, coordinate-independent variance built from DeWitt-type momentum operators. In this sense, the familiar single-slit thought experiment is lifted from flat space to arbitrary slices of a Lorentzian spacetime.
From this viewpoint, the main intrinsic estimate can be read as a spectral version of Heisenberg’s principle. For any normalized state $\psi$, strict localization in the ball enforces the lower bound
$\sigma_{p}(\psi)\;\geq\;\hbar\,\sqrt{\lambda_{1}}\,$,
where $\lambda_{1}$ is the first Dirichlet eigenvalue of the Laplace–Beltrami operator on the ball. The bound is intrinsic in a strong sense: lapse, shift and extrinsic curvature never enter, and only the induced Riemannian metric on the slice survives. Physically, $\lambda_{1}$ condenses into a single number all geometric features of the localization region that matter for the trade-off between position and momentum: its proper radius and volume, the shape and mean curvature of its boundary, and the curvature profile in the interior. The kinetic-energy window separating the gradients of $|\psi|$ and of $\psi$ itself shows that a nontrivial part of the momentum uncertainty is already fixed by the modulus of the wave function, while fluctuations of the phase gradient only add a nonnegative excess. The equality conditions make it clear that the lower bound is sharp but typically not saturated on curved balls, reflecting the fact that a globally constant phase gradient is hard to realize in a nontrivial geometry.
In many situations one does not know $\lambda_{1}$ explicitly, but one does have qualitative control over the boundary of the ball. Under the weak mean-convexity assumption, the boundary-distance Hardy inequality and the vector-field version of Barta’s method lead to universal, scale-invariant product bounds
$\sigma_{p}\,r\;\geq\;\frac{\hbar}{2}\,,\qquad\sigma_{p}\,r\;\geq\;\frac{\pi\hbar}{2}\,$,
for all strictly localized states. These inequalities have several noteworthy features. They depend only on the proper radius of the ball and the sign of the radial Laplacian; the detailed curvature in the interior is irrelevant. The Hardy baseline with constant $1/2$ is never attained, in line with the flat-space Hardy inequality, which underscores the robustness of the bound against small perturbations of the state. The Barta-type improvement effectively extracts everything that can be said from mean-convexity alone; any further sharpening of the constant must bring in additional curvature information, for instance via classical eigenvalue comparison theorems or explicit model calculations on constant-curvature spaces.
The examples in Section 4 illustrate how these general statements play out in concrete geometries. On unimodular group manifolds such as the three-dimensional Heisenberg (Nil) geometry, the geometric potential $V$ vanishes even though the scalar curvature is negative, so the intrinsic momentum uncertainty is governed entirely by the Dirichlet spectrum on geodesic balls. On hyperbolic space $H^{3}$ the non-unimodular Lie algebra generates a positive constant $V$ that coincides with the spectral gap of the Dirichlet Laplacian on large balls. In this case the intrinsic variance separates neatly into a confinement contribution with the familiar $1/r^{2}$ scaling and a curvature-induced offset. Witten’s cigar provides an example with non-constant curvature: here $V(r)$ is a genuine potential that regularizes the effective Hamiltonian near the tip and decays along the asymptotically cylindrical region. Finally, the maximum-length model from the GUP literature shows that a purely algebraic deformation of the commutator can be reinterpreted as standard quantum mechanics on a conformally flat configuration space. In that rereading, the intrinsic lower bound on $\sigma_{p}$ is dictated by the Dirichlet spectrum of the reconstructed metric, and the deformation parameter plays the role of a curvature scale.
Conceptually, these observations suggest a different perspective on generalized or extended uncertainty relations. Instead of taking them as a fundamental departure from the canonical commutation relations, one may regard them as effective manifestations of the spectral geometry of the localization regions available in a given gravitational setting. The present framework demonstrates that on any spacelike hypersurface the interplay between localization and momentum can be quantified without leaving standard quantum mechanics on curved spacetime: once the induced Riemannian metric and a localization region are specified, the spectral geometry of that region fixes a lower bound on the momentum standard deviation $\sigma_{p}$. In this sense, the “minimal lengths” and “maximal localization scales” that often appear in phenomenological GUP models need not be new constants of nature; they can emerge as geometric scales such as the injectivity radius, curvature radius or throat size of the slice.
At the same time, the limitations of the present treatment point toward several natural extensions. Throughout we worked with test particles on a fixed background, so that backreaction of the localized quantum state on the geometry was ignored. In regimes where the momentum spread demanded by the intrinsic bounds approaches the Planck scale, one expects significant gravitational backreaction and possibly horizon formation; a self-consistent treatment would then have to determine the spectrum of $B_{\Sigma}(p,r)$ jointly with the Einstein equations. A second direction is the generalization from single-particle quantum mechanics to quantum field theory on curved backgrounds. There the notion of localization is more subtle, but the spectral point of view suggests that analogous bounds might be formulated for suitably smeared field operators or for local algebras associated with geodesic balls. Finally, in the context of quantum sensing, the universal product bounds derived here can be interpreted as hard limits on the spatial resolution of probes that are confined to strongly curved regions such as black-hole throats or cosmological horizons, and may serve as design constraints for measurement schemes that try to balance spatial focus against momentum or energy noise.
To summarize, the results obtained in this work recast Heisenberg’s uncertainty principle as a statement about the intrinsic spectral geometry of spacelike hypersurfaces in general relativity. Strict localization to a geodesic ball carries an unavoidable spectral cost, quantified by the first Dirichlet eigenvalue and, under weak convexity assumptions, bounded below by universal constants that are independent of the foliation and insensitive to the detailed interior curvature. These intrinsic Heisenberg-type inequalities show that quantum uncertainty and spacetime geometry are intertwined already at the level of single-particle quantum mechanics on curved backgrounds, well before any explicit quantization of the gravitational field is invoked.
Ultimately, this implies that the uncertainty principle ceases to be merely a kinematic restriction of the state vector, but reveals itself as a geometric necessity enforced by the curvature of spacetime itself.